A modern-day zeppelin holds 7,490 m3 of helium. Compute its maximum payload at sea level. (Assume the helium and air to be at 0°C and 1 atm.)
1
Expert's answer
2019-04-15T10:08:12-0400

The buoyant force is


"F_b=\rho_{air} g V"

The weight of the helium:

"W=\rho_{He} g V"

The maximum payload at sea level:


"F=F_b-W=(\rho_{air}-\rho_{He}) g V"

"F=(1.22-0.18)(9.81)(7490)=76400N=76.4 kN"